# SurfPro: Functional Protein Design Based on Continuous Surface

### v1

@@ 4.2 Inverse Folding
Table 2: AF2 pAE_interaction ($\downarrow$) for all models in the binder design task. “Average” denotes the average AF2 pAE_interaction across the entire test set instead of the direct average on different target proteins. We also provide the AF2 pAE_interaction for randomly sampled negative binders of the same length as the positive ones.
@@
Table 3: Success rate (%, $\uparrow$) for all models in the binder design task. “Average” denotes the average success rate across the entire test set instead of the direct average on different target proteins.
@@ 4.3 Protein Binder Design
Table 5: Success rate (%, $\uparrow$) for all models in the enzyme design task. “Average” denotes the average success rate across the entire test set instead of the direct average on different enzyme categories. The substrate is denoted as its KEGG database ID.

### v2

@@ 4.2 Inverse Folding
Table 2: AF2 pAE_interaction ($\downarrow$) for all models in the binder design task. “Average” denotes the average AF2 pAE_interaction across the entire test set instead of the direct average on different target proteins. We also provide the AF2 pAE_interaction for randomly sampled negative binders of the same length as the positive ones. Our SurfPro outperforms all previous methods on AF2 pAE_interaction.
@@
Table 3: Success rate (%, $\uparrow$) for all models in the binder design task. “Average” denotes the average success rate across the entire test set instead of the direct average on different target proteins. Our SurfPro-Pretrain outperforms all previous methods in success rate by a big margin.
@@ 4.3 Protein Binder Design
Table 5: Success rate (%, $\uparrow$) for all models in the enzyme design task. “Average” denotes the average success rate across the entire test set instead of the direct average on different enzyme categories. The substrate is denoted as its KEGG database ID. Notice that our SurfPro achieves the highest average success rate.
@@ 5.1 Ablation Study: How Does Each Component Work?
Methods Perplexity ($\downarrow$) Recovery Rate (%, $\uparrow$)
SurfPro 3.13 57.78
– w NAD 12.51 9.61
Table 7: Ablation study on CATH 4.2 dataset. ($\uparrow$): the higher the better. ($\downarrow$): the lower the better. NAD denotes non-autoregressive decoder.
5.2 Ablation Study: Can Autoregressive Decoder Be Replaced?
The autoregressive decoder demonstrates strong performance within our proposed SurfPro architecture. To assess its significance, we compare SurfPro with an alternative employing a non-autoregressive decoder, incorporating SoftCopy and glancing learning strategies inspired by non-autoregressive machine translation in natural language processing (Gu & Kong, 2021; Qian et al., 2021). Table 7 presents the comparative results. It unequivocally demonstrates that SurfPro outperforms the non-autoregressive decoder variant, affirming the efficacy of the autoregressive decoder within our proposed framework.
@@ 5.4 How Does Vertex Number Affect Protein Design?
Models Seen Class Zero-Shot Average Inference Time (s/sample)
InsulinR PDGFR TGFb H3 IL7Ra TrkA
MaSIF 9.0737 10.8734 18.8502 21.9093 25.9689 25.9112 16.1280 210.42
SurfPro 10.2608 17.9862 17.7364 21.2916 20.8594 10.6535 16.9485 0.45
Table 8: Comparison with MaSIF on binder design task. “Average” denotes the average AF2 pAE_interaction across the entire test set instead of the direct average on different target proteins.
5.5 Comparing with MaSIF
In this section, we conduct a comparative analysis between our SurfPro and the well-established surface-based binder design model MaSIF (Gainza et al., 2020). Initially, we employ ProteinMPNN to generate 100 binder sequences for each positive binder. Subsequently, ESMFold is utilized to predict the structures of these designed binder sequences. Following this, MaSIF is employed to rank these candidate binders, with the top-performing one selected as the final designed binder. The resulting pAE_interaction scores across six datasets are summarized in Table 8. The comparison reveals that MaSIF outperforms our model in 2 out of 6 target proteins, while our SurfPro outperforms in 4 out of 6 target proteins. Despite MaSIF demonstrating a slightly superior average pAE_interaction score, it’s noteworthy that the average time taken for MaSIF to identify a promising binder is 210.42 seconds, significantly longer than the mere 0.45 seconds required by SurfPro. This underscores SurfPro’s efficacy as a standout generative model, capable of swiftly and directly generating functional protein sequences in an end-to-end manner.
@@ 5.6 Case Study
6 Discussion
Our SurfPro exhibits superior performance in rapidly and directly generating functional protein sequences based on provided protein surfaces. However, despite its impressive capabilities, certain limitations persist, which we will discuss in this section. While our approach excels in protein optimization, it leans more towards refinement rather than de novo protein design. This distinction is significant and meaningful. Particularly in binder design, achieving a high-affinity binder from scratch is seldom feasible in practical scenarios. Thus, starting from existing positive binders can accelerate the design process. Nonetheless, this approach also imposes constraints on the practical utility of our method. Firstly, locating a positive initial point isn’t always feasible. Secondly, commencing from a favorable starting point may result in limited improvements, as our SurfPro necessitates consideration of both geometric and biochemical constraints. Our future work could explore methods for de novo designing protein surfaces. For example, integrating diffusion models to generate point clouds could substantially enhance the versatility and applicability of our existing framework.
@@ 7 Conclusion
Acknowledgement
This research is partially supported by the UC Santa Barbara Faculty Research Grant (to L.L.). The authors thank the anonymous reviewers and Wenxian Shi, Siqi Ouyang, Danqing Wang, and Yujia Gao for their excellent suggestions.